You would like to deposit $1000 in a savings account. The bank gives you two options: Account A pays interest at a rate of 8% (annually compounded) each year. Account B pays a simple interest rate of 8% per year (but never compounds), and gives you a $50 bonus at the end of 5 years. What is the difference in the account balances in 5 years? (Your answer should be the solution to Balance of Account A minus Balance of Account B.)
Answer #1
Balance in Account A $ 1,469.33
Minus Balance in Account B $ 1,450.00
Difference in the account balances in 5 years $       19.33
Working:
# 1 Account balance in 5 years = P*(1+i)^n Where,
(Option A) = 1000*(1+0.08)^5 P $       1,000
= $ 1,469.33 i 8%
n 5
# 2 Account balance in 5 years = Investment + Interest Earned + Bonus
(Option B) = 1000+(1000*8%*5)+50
= $ 1,450.00